# Decomposable Submodular Maximization in Federated Setting

### v1

@@ 1 Introduction
Decomposable submodular functions include some of the most fundamental and well-studied submodular functions such as max coverage, graph cuts, welfare maximization etc., and have found numerous applications in machine learning (Dueck and Frey, 2007; Gomes and Krause, 2010; Mirzasoleiman, Badanidiyuru, and Karbasi, 2016a, b; Mirzasoleiman, 2017), economics (Dobzinski and Schapira, 2006; Feige, 2006; Feige and Vondrák, 2006; Papadimitriou et al., 2008; Vondrák, 2008), social network (Kempe et al., 2003), and data summarization and recommender systems (Dueck and Frey, 2007; Gomes and Krause, 2010; Tschiatschek et al., 2014; Lin and Bilmes, 2011).
As a consequence of these wide-prevailing applications, maximizing decomposable submodular functions subject to different constraints has attracted significant theoretical/practical interest. The main approach in these cases is the centralized and sequential greedy approach.
@@
The welfare maximization problem aims to maximize the overall utility or welfare of a group of individuals or agents $a_{1},\dots,a_{N}$. In this problem, there is a set of items or goods $E$, and each individual $a_{i}$ has a certain preference or utility expressed as a submodular function $f_{i}:2^{E}\to\mathbb{R}_{+}$ that assigns a value to each combination of items. The goal is to partition $E$ into disjoint subsets $S_{1},\dots,S_{N}$ in order to maximize the social welfare $\sum_{i=1}^{N}f_{i}(S_{i})$. In an important special case, called Combinatorial Public Projects (Gupta et al., 2010; Papadimitriou et al., 2008), the goal is to find a subset $S\subseteq E$ of size at most $k$ that maximizes $F(S)=\sum_{i=1}^{N}f_{i}(S)$. This problem can be applied in different fields, such as resource allocation, public goods provision, market design, and has been intensively studied (Khot et al., 2008; Lehmann et al., 2006; Mirrokni et al., 2008). An optimal approximation algorithm is known for this problem in value oracle model in which it is required to have access to the value of $f_{i}(S)$ for each agent and any $S\subseteq E$ (Călinescu et al., 2011). However, in scenarios where agents are hesitant to disclose their data to a central server and storing all data on a single server is computationally infeasible, the demand for a decentralized and private submodular maximization algorithm becomes imperative.
@@
Enabling privacy-protected data sharing among clinical centers is crucial for global collaborations. In the context of feature selection over sensitive medical data, consider geographically dispersed hospitals, denoted as $H_{1},\dots,H_{N}$, with each hospital maintaining its own data that it is unwilling to share. The goal is to identify a small subset $S$ of features that effectively classifies the target variable across the entire dataset over all hospitals. The demand for a decentralized and private feature selection algorithm is paramount to discover hidden patterns while preserving data ownership. By adopting a decentralized approach, hospitals can balance collaborative knowledge discovery and data privacy. One approach is to maximize a submodular function capturing the mutual information between features and the class labels (Krause and Guestrin, 2005).
@@
Problem definition and setting. In this paper we introduce the problem of maximizing a submodular function in the federated setting. Let $E$ be a ground set of size $n$ and $c_{1},\dots,c_{N}$ be $N$ clients each of whom has a private interest over $E$. Each client’s interest is expressed as a submodular function. Let $f_{i}:2^{E}\to\mathbb{R}_{+}$ be the associated submodular function of the $i$-th client. A central server wants to solve the following constrained distributed optimization model
@@
where $\mathcal{I}$ is the independent sets of a matroid $\mathcal{M}$ with ground set $E$, and $p_{i}$ are pre-defined weights such that $\sum\limits_{i=1}^{N}p_{i}=1$. For instance, they can be set to $1/N$, or the fraction of data owned by each client. The constraint implies sets of particular properties, e.g., subsets of size at most $k$. Note that, the unconstrained optimization is a special case of this.
In the optimization problem (1) the data can be massively distributed over the number of clients $N$, which can be huge. Moreover unlike the traditional distributed setting, in the federated setting the server does not have control over clients’ devices nor on how data is distributed. For example, when a mobile phone is turned off or WiFi access is unavailable, the central server will lose connection to this device. Furthermore, client’s objective can be very different depending on their local datasets. To minimize communication overhead and server computation load, the number of communication rounds from clients to servers need to be minimal.
@@
  • 1.
    Privacy: One of the main appeals of decentralized and federated setting is to preserve the privacy. There are several models of privacy and security that have been considered in the literature such as Differential Privacy (DP) and Secure Aggregator (SecAgg), and a mix of these two. Single-server SecAgg is a cryptographic secure multi-party computation (MPC) that enables clients to submit vector inputs, such that the server (an aggregator) can only decipher the combined update, not individual updates. This is usually achieved via additive masking over a finite group (Bell et al., 2020; Bonawitz et al., 2016). Note that secure aggregation alone does not provide any privacy guarantees. To achieve a DP-type guarantee, noise can be added locally, with the server aggregating the perturbed local information via SecAgg. This user-level DP framework has recently been adopted in private federated learning (Agarwal et al., 2018, 2021; Kairouz et al., 2021; Chen et al., 2022a).
@@
  • 2.
    Communication and bit complexity: There are a few aspects to this, firstly the number of communication rounds between the server and clients should be as small as possible. Second, the information communicated between should require low bandwidth and they better require small bit complexity to encode.
  • 3.
    Convergence and utility: While the above impose strong restrictions, a good decentralized submodular maximization algorithm should not scarify the convergence rate by too much and should yield to an accurate and acceptable result in comparison to the centeralized methods.
Our contributions.
We present the first federated (constrained) submodular maximization algorithm that converges close to optimum guarantees known in centralized settings.
$\bullet$ We propose a decentralized version of the popular Continuous Greedy algorithm Federated Continuous Greedy (FedCG) and prove its convergence whenever the client functions are nonnegative monotone submodular achieving the optimal multiplicative approximation factor $(1-\nicefrac{{1}}{{e}})$ with a small additive.
$\bullet$ We incorporate important and practical scenarios that are relevant for federated setting such as partial client selection, low communication rounds and computation cost. We give rigorous theoretical guarantees under each scenarios matching the optimal multiplicative approximation of $(1-\nicefrac{{1}}{{e}})$ and small additive error.
@@ 1.1 Related work
The concept of Federated Learning (FL), introduced by McMahan et al. (2017), has found application in various domains such as natural language processing, computer vision, and healthcare. The popular FL algorithm, Federated Averaging (FedAvg), is an extension of Local SGD that aims to reduce communication costs in distributed settings (Gorbunov et al., 2021; Stich, 2019; Wang and Joshi, 2021; Yu et al., 2019). However, despite its practical benefits in addressing privacy, data heterogeneity, and computational constraints, it may not converge to a “good enough” solution in general (Pathak and Wainwright, 2020; Zhang et al., 2020), and analyzing the convergence of FedAvg and providing theoretical guarantees is challenging and necessitates making certain assumptions. Assumptions related to bounded gradients, convexity, Lipschitzness, statistical heterogeneity, and bounded variance of stochastic gradients for each client have been explored in recent works (Karimireddy et al., 2020; Li et al., 2020; Stich, 2019; Woodworth et al., 2020; Wang et al., 2019; Yu et al., 2019).
@@
Decentralized / distributed submodular maximization. The main approach to submodular maximization is the greedy approach which in fact, in the centralized setting yields the tight approximation guarantee in various scenarios and constraints e.g., see (Nemhauser et al., 1978; Vondrák, 2008; Călinescu et al., 2011). Centeralized submodular maximization under privacy constraints is an active research area (Mitrovic et al., 2017; Rafiey and Yoshida, 2020; Chaturvedi et al., 2021). However the sequential nature of the greedy approach makes it challenging to scale it to massive datasets. This issue is partially addressed by the means of Map-Reduce style algorithms (Kumar et al., 2015) as well as several elegant algorithms in the distributed setting (Mirzasoleiman et al., 2016b; Barbosa et al., 2015). Recent work of Mokhtari et al. (2018b) ventures towards decentralized submodular maximization for continuous submodular functions. Note that in general continuous submodular functions are not convex nor concave and there has been a line of work to optimize continuous submodular functions using SGD methods (Hassani et al., 2017). Mokhtari et al. under several assumptions, such as assuming clients’ local objective functions are monotone, DR-submodular, Lipschitz continuous, and have bounded gradient norms, prove that Decenteralized Continuous Greedy algorithm yields a feasible solution with quality $O(1-1/\mathrm{e})$ times the optimal solution. The setting in (Mokhtari et al., 2018b) is fundamentally different from the federated setting in a sense that they require sharing gradient information of the clients with the server or with the neighboring nodes in an underlying graph. Perhaps the most closely related method to our work is due to Dadras et al. (2022); Zhang et al. (2022). Dadras et al. (2022) consider Frank-Wolfe Algorithm (Frank and Wolfe, 1956) in the federated setting and propose Federated Frank-Wolfe (FedFW) algorithm and analyze its convergence for both convex and non-convex functions and under the $L$-Lipschitzness and bounded gradients assumptions.
@@ 2 Preliminaries
Let $E$ denote the ground set and $|E|=n$. For a vector $\mathbf{x}\in\mathbb{R}^{|E|}$ and a set $S\subseteq E$, we use $\mathbf{x}(S)$ to denote $\sum_{e\in S}\mathbf{x}(e)$. We say a submodular function $f:2^{E}\to\mathbb{R}$ is monotone if $f(S)\leq f(T)$ for every $S\subseteq T\subseteq E$. Throughout this paper we assume $f(\emptyset)=0$.
@@
where $R$ is the random subset of $E$ in which each element $e\in E$ is included independently with probability $\mathbf{x}(e)$ and is not included with probability $1-\mathbf{x}(e)$. We write $R\sim\mathbf{x}$ to denote that $R\subseteq E$ is a random set sampled according to $\mathbf{x}$.
Observe that for all $S\subseteq E$ we have $\widehat{f}(\mathbf{1}_{S})=f(S)$. For monotone non-decreasing submodular function $f$, $\widehat{f}$ has the following properties (Călinescu et al., 2011) that are crucial in analyses of our algorithms
  • 1.
    $\widehat{f}$ is monotone, meaning $\frac{\partial\widehat{f}}{\partial\mathbf{x}(e)}\geq 0$. Hence, $\nabla\widehat{f}(\mathbf{x})=(\frac{\partial\widehat{f}}{\partial\mathbf{x}(1)},\dots,\frac{\partial\widehat{f}}{\partial\mathbf{x}(n)})$ is a nonnegative vector.
  • 2.
    $\widehat{f}$ is concave along any direction $\mathbf{d}\geq\mathbf{0}$.
@@
Matroids and matroid polytopes.
A pair $\mathcal{M}=(E,\mathcal{I})$ of a set $E$ and $\mathcal{I}\subseteq 2^{E}$ is called a matroid if 1) $\emptyset\in\mathcal{I}$, 2) $A\in\mathcal{I}$ for any $A\subseteq B\in\mathcal{I}$, and 3) for any $A,B\in\mathcal{I}$ with $|A|<|B|$, there exists $e\in B\setminus A$ such that $A\cup\{e\}\in\mathcal{I}$. We call a set in $\mathcal{I}$ an independent set. We sometimes abuse notation and use $S\in\mathcal{M}$. The rank function $r_{\mathcal{M}}\colon 2^{E}\to\mathbb{Z}_{+}$ of $\mathcal{M}$ is $r_{\mathcal{M}}(S)=\max\{|I|:I\subseteq S,I\in\mathcal{I}\}$. An independent set $S\in\mathcal{I}$ is called a base if $r_{\mathcal{M}}(S)=r_{\mathcal{M}}(E)$. We denote the rank of $\mathcal{M}$ by $r(\mathcal{M})$. The matroid polytope $\mathcal{P}(\mathcal{M})\subseteq\mathbb{R}^{E}$ of $\mathcal{M}$ is $\mathcal{P}(\mathcal{M})=\mathrm{conv}\{\mathbf{1}_{I}:I\in\mathcal{I}\}$ where $\mathrm{conv}$ denotes the convex hull. Or equivalently (Edmonds, 2001), $\mathcal{P}(\mathcal{M})=\left\{\mathbf{x}\geq\mathbf{0}:\mathbf{x}(S)\leq r_{\mathcal{M}}(S);~{}\forall S\subseteq E\right\}.$ The base polytope of $\mathcal{M}$ is given by $B(\mathcal{M})=\{\mathbf{y}\in\mathcal{P}(\mathcal{M})\mid\mathbf{y}(E)=r_{\mathcal{M}}(E)\}$. The extreme points of $B(\mathcal{M})$ are the characteristic vectors of the bases of $\mathcal{M}$.
The Continuous Greedy Algorithm.
Our algorithms for maximizing a submodular function in federated settings are based on the Continuous Greedy (CG) algorithm. Here we briefly explain this algorithm. The results mentioned are from (Călinescu et al., 2011; Vondrák, 2008). Let $\mathcal{M}=(E,\mathcal{I})$ be a matroid and $\mathcal{P}(\mathcal{M})$ be its matroid polytope of rank $r$, let $f$ be a nonnegative and monotone submodular function and $\widehat{f}$ be its multilinear extension. CG starts with $\mathbf{x}^{(0)}=\mathbf{0}$. For every $t\in\{0,1,2,\dots,T-1\}$ it computes $\mathbf{x}^{(t+1)}$ using the following update step $\mathbf{x}^{(t+1)}\leftarrow\mathbf{x}^{(t)}+\eta\mathbf{v}^{(t)}$, where $\mathbf{v}^{(t)}=\operatornamewithlimits{argmax}_{\mathbf{w}\in\mathcal{P}}\langle\mathbf{w},\nabla\widehat{f}(\mathbf{x}^{(t)})\rangle$. For $\texttt{OPT}=\max_{\mathbf{x}\in\mathcal{P}}\widehat{f}(\mathbf{x})$ we have $(1-(1-\eta)^{T})\texttt{OPT}\leq\widehat{f}(\mathbf{x}^{(T)})+C\eta^{2}/2$. Here, the constant $C$ depends on the Lipschitz of the function, and $\mathbf{x}^{(T)}\in\mathcal{P}$ as it is a convex combination of vectors from the polytope. Moreover, for $\eta=1/T$ and large enough $T$ we get $(1-1/\mathrm{e})\texttt{OPT}\leq\widehat{f}(\mathbf{x}^{(T)})+\epsilon$.
Given $\mathbf{x}^{(T)}\in\mathcal{P}$, there are rounding procedures to obtain $S\in\mathcal{I}$ such that $\widehat{f}(\mathbf{x}^{(T)})\leq f(S)$. The approximation factor $1-1/\mathrm{e}$ is the best possible assuming $\mathbf{P}\neq\mathbf{NP}$ (Feige, 1998).

### v2

@@ 1 Introduction
Decomposable submodular functions include some of the most fundamental and well-studied submodular functions such as max coverage, graph cuts, welfare maximization etc., and have found numerous applications in machine learning (Dueck and Frey, 2007; Gomes and Krause, 2010; Mirzasoleiman, Badanidiyuru, and Karbasi, 2016a, b; Mirzasoleiman, 2017), economics (Dobzinski & Schapira, 2006; Feige, 2006; Feige & Vondrák, 2006; Papadimitriou et al., 2008; Vondrák, 2008), and data summarization and recommender systems (Dueck & Frey, 2007; Gomes & Krause, 2010; Tschiatschek et al., 2014; Lin & Bilmes, 2011). The main approach in these cases is the centralized and sequential greedy.
@@
The welfare maximization problem aims to maximize the overall utility or welfare of a group of individuals or agents $a_{1},\dots,a_{N}$. In this problem, there is a set of items or goods $E$, and each individual $a_{i}$ has a certain preference or utility expressed as a submodular function $f_{i}:2^{E}\to\mathbb{R}_{+}$ that assigns a value to each combination of items. The goal is to partition $E$ into disjoint subsets $S_{1},\dots,S_{N}$ in order to maximize the social welfare $\sum_{i=1}^{N}f_{i}(S_{i})$. In an important special case, called Combinatorial Public Projects (Gupta et al., 2010; Papadimitriou et al., 2008), the goal is to find a subset $S\subseteq E$ of size at most $k$ maximizing $F(S)=\sum_{i=1}^{N}f_{i}(S)$. This problem appears in different fields, such as resource allocation, public goods provision, market design, and has been intensively studied (Khot et al., 2008; Lehmann et al., 2006; Mirrokni et al., 2008). An optimal approximation algorithm is known in value oracle model in which it is required to have access to the value of $f_{i}(S)$ for each agent and any $S\subseteq E$ (Călinescu et al., 2011). However, in scenarios where agents are hesitant to disclose their data to a central server and storing all data on a single server is computationally infeasible, the demand for a decentralized and private submodular maximization algorithm becomes imperative.
@@
Enabling privacy-protected data sharing among clinical centers is crucial for global collaborations. Consider geographically dispersed hospitals $H_{1},\dots,H_{N}$, with each hospital maintaining its own data that it is unwilling to share. The goal is to identify a small subset of features that effectively classifies the target variable across the entire dataset over all hospitals. A decentralized and efficient feature selection algorithm is crucial for uncovering hidden patterns while maintaining data ownership. By adopting a decentralized approach, hospitals can balance collaborative knowledge discovery and data privacy. One approach is to maximize a submodular function capturing the mutual information between features and the class labels (Krause & Guestrin, 2005).
@@
Problem definition and setting. In this paper, we introduce the problem of maximizing a submodular function in the federated setting. Let $E$ be a ground set of size $n$ and $c_{1},\dots,c_{N}$ be $N$ clients each of whom has a private interest over $E$. Each client’s interest is expressed as a submodular function. Let $f_{i}:2^{E}\to\mathbb{R}_{+}$ be the associated submodular function of the $i$-th client. A central server wants to solve the following constrained distributed optimization model
@@
where $\mathcal{I}$ is the independent sets of a matroid $\mathcal{M}$ with ground set $E$, and $p_{i}$ are pre-defined weights such that $\sum_{i=1}^{N}p_{i}=1$. For instance, they can be set to $1/N$, or the fraction of data owned by each client. The constraint implies sets of particular properties, e.g., subsets of size at most $k$. Note that, the unconstrained optimization is a special case of this.
In the optimization problem (1) the data can be massively distributed over the number of clients $N$, which can be huge. Moreover unlike the traditional distributed setting, in the federated setting the server does not have control over clients’ devices nor on how data is distributed. For example, when a mobile phone is turned off or WiFi access is unavailable, the central server will lose connection to this device. Furthermore, client’s objective can be very different depending on their local datasets. To minimize communication overhead and server computation load, the number of communication rounds need to be minimal.
@@
1. Privacy: One of the main appeals of decentralized and federated setting is to preserve the privacy. There are several models of privacy and security that have been considered in the literature such as Differential Privacy (DP) and Secure Aggregator (SecAgg), and a mix of these two. Single-server SecAgg is a cryptographic secure multi-party computation (MPC) that enables clients to submit vector inputs, such that the server (an aggregator) can only decipher the combined update, not individual updates. This is usually achieved via additive masking over a finite group (Bell et al., 2020; Bonawitz et al., 2016). Note that secure aggregation alone does not provide any privacy guarantees. To achieve a DP-type guarantee, noise can be added locally, with the server aggregating the perturbed local information via SecAgg. This user-level DP framework has recently been adopted in private federated learning (Agarwal et al., 2018, 2021; Kairouz et al., 2021; Chen et al., 2022a; Wang et al., 2023).
@@
2. Communication and bit complexity: There are a few aspects to this, firstly the number of communication rounds should be as small as possible. Second, the information communicated between should require low bandwidth and they better require small bit complexity to encode.
3. Convergence and utility: While the above impose strong restrictions, a good decentralized submodular maximization algorithm should not scarify the convergence rate by too much and should yield to an accurate and acceptable result in comparison to the centeralized methods.
Our contributions. We present the first federated (constrained) submodular maximization algorithm converging close to optimum guarantees known in centralized settings.
$\bullet$ We propose a decentralized version of the popular Continuous Greedy algorithm Federated Continuous Greedy (FedCG) and prove its convergence whenever the client functions are nonnegative monotone submodular achieving the optimal multiplicative approximation factor $(1-\nicefrac{{1}}{{e}})$ with a small additive (Section 3).
$\bullet$ We incorporate important and practical scenarios that are relevant for federated setting such as partial client selection, low communication rounds and computation cost. We give rigorous theoretical guarantees under each scenarios matching the optimal multiplicative approximation of $(1-\nicefrac{{1}}{{e}})$ and small additive error (Section 4).
@@ 1.1 Related Work
The concept of Federated Learning (FL) (McMahan et al., 2017) has found application in various domains such as natural language processing, computer vision, and healthcare. The popular FL algorithm, Federated Averaging (FedAvg), is an extension of Local SGD that aims to reduce communication costs in distributed settings (Gorbunov et al., 2021; Stich, 2019; Wang & Joshi, 2021; Yu et al., 2019). However, despite its practical benefits in addressing privacy, data heterogeneity, and computational constraints, it may not converge to a “good enough” solution in general (Pathak & Wainwright, 2020; Zhang et al., 2020). Analyzing the convergence of FedAvg and providing theoretical guarantees is challenging and necessitates making certain assumptions. Assumptions related to bounded gradients, convexity, Lipschitzness, statistical heterogeneity, and bounded variance of stochastic gradients for each client have been explored in recent works (Karimireddy et al., 2020; Li et al., 2020; Woodworth et al., 2020; Wang et al., 2019; Yu et al., 2019).
@@
Decentralized / distributed submodular maximization. The main approach to submodular maximization is the greedy approach which in fact, in the centralized setting yields the tight approximation guarantee in various scenarios and constraints e.g., see (Nemhauser et al., 1978; Vondrák, 2008; Călinescu et al., 2011). Centeralized submodular maximization under privacy constraints is an active research area (Mitrovic et al., 2017; Rafiey & Yoshida, 2020; Chaturvedi et al., 2021). However the sequential nature of the greedy approach makes it challenging to scale it to massive datasets. This issue is partially addressed by the means of Map-Reduce style algorithms (Kumar et al., 2015) as well as several elegant algorithms in the distributed setting (Mirzasoleiman et al., 2016b; Barbosa et al., 2015). Recent work of Mokhtari et al. (2018b) ventures towards decentralized submodular maximization for continuous submodular functions. In general continuous submodular functions are not convex nor concave and there has been a line of work to optimize continuous submodular functions using SGD methods (Hassani et al., 2017). Mokhtari et al. under several assumptions, such as assuming clients’ local objective functions are monotone, DR-submodular, Lipschitz continuous, and have bounded gradient norms, prove that Decenteralized Continuous Greedy algorithm yields a feasible solution with quality $O(1-1/\mathrm{e})$ times the optimal solution. The setting in (Mokhtari et al., 2018b) is fundamentally different from the federated setting in a sense that they require sharing gradient information of the clients with the server or with the neighboring nodes in an underlying graph. Perhaps the most closely related method to our work is due to Dadras et al. (2022); Zhang et al. (2022). Dadras et al. (2022) consider Frank-Wolfe Algorithm (Frank & Wolfe, 1956) in the federated setting and propose Federated Frank-Wolfe (FedFW) algorithm and analyze its convergence for both convex and non-convex functions and under the $L$-Lipschitzness and bounded gradients assumptions.
@@ 2 Preliminaries
Let $E$ denote the ground set, $|E|=n$. For a vector $\mathbf{x}\in\mathbb{R}^{|E|}$ and a set $S\subseteq E$, $\mathbf{x}(S)$ denotes $\sum_{e\in S}\mathbf{x}(e)$. A submodular function $f:2^{E}\to\mathbb{R}$ is monotone if $f(S)\leq f(T)$ for every $S\subseteq T\subseteq E$. Throughout this paper we assume $f(\emptyset)=0$.
@@
where $R\subseteq E$ is a random set that contains each element $e\in E$ independently with probability $\mathbf{x}(e)$ and excludes it with probability $1-\mathbf{x}(e)$. We write $R\sim\mathbf{x}$ to denote that $R\subseteq E$ is a random set sampled according to $\mathbf{x}$.
Observe that for all $S\subseteq E$ we have $\widehat{f}(\mathbf{1}_{S})=f(S)$. For monotone non-decreasing submodular function $f$, $\widehat{f}$ has the following properties (Călinescu et al., 2011) that are crucial in analyses of our algorithms:
1. $\widehat{f}$ is monotone, meaning $\frac{\partial\widehat{f}}{\partial\mathbf{x}(e)}\geq 0$. Hence, $\nabla\widehat{f}(\mathbf{x})=(\frac{\partial\widehat{f}}{\partial\mathbf{x}(1)},\dots,\frac{\partial\widehat{f}}{\partial\mathbf{x}(n)})$ is a nonnegative vector.
2. $\widehat{f}$ is concave along any direction $\mathbf{d}\geq\mathbf{0}$.
@@
Matroids and matroid polytopes. A pair $\mathcal{M}=(E,\mathcal{I})$ of a set $E$ and $\mathcal{I}\subseteq 2^{E}$ is called a matroid if 1) $\emptyset\in\mathcal{I}$, 2) $A\in\mathcal{I}$ for any $A\subseteq B\in\mathcal{I}$, and 3) for any $A,B\in\mathcal{I}$ with $|A|<|B|$, there exists $e\in B\setminus A$ such that $A\cup\{e\}\in\mathcal{I}$. We call a set in $\mathcal{I}$ an independent set. We sometimes abuse notation and use $S\in\mathcal{M}$. The rank function $r_{\mathcal{M}}\colon 2^{E}\to\mathbb{Z}_{+}$ of $\mathcal{M}$ is $r_{\mathcal{M}}(S)=\max\{|I|:I\subseteq S,I\in\mathcal{I}\}$. An independent set $S\in\mathcal{I}$ is called a base if $r_{\mathcal{M}}(S)=r_{\mathcal{M}}(E)$. We denote the rank of $\mathcal{M}$ by $r(\mathcal{M})$. The matroid polytope $\mathcal{P}(\mathcal{M})\subseteq\mathbb{R}^{E}$ of $\mathcal{M}$ is $\mathcal{P}(\mathcal{M})=\mathrm{conv}\{\mathbf{1}_{I}:I\in\mathcal{I}\}$ where $\mathrm{conv}$ denotes the convex hull. Or equivalently (Edmonds, 2001), $\mathcal{P}(\mathcal{M})=\left\{\mathbf{x}\geq\mathbf{0}:\mathbf{x}(S)\leq r_{\mathcal{M}}(S);~{}\forall S\subseteq E\right\}.$
The Continuous Greedy Algorithm. Our algorithms for maximizing a submodular function in federated settings are based on the Continuous Greedy (CG) algorithm. We briefly explain this algorithm. The results mentioned are from (Călinescu et al., 2011; Vondrák, 2008). Let $\mathcal{M}=(E,\mathcal{I})$ be a matroid and $\mathcal{P}(\mathcal{M})$ be its matroid polytope of rank $r$, let $f$ be a nonnegative and monotone submodular function and $\widehat{f}$ be its multilinear extension. CG starts with $\mathbf{x}^{(0)}=\mathbf{0}$. For every $t\in\{0,1,2,\dots,T-1\}$ it computes $\mathbf{x}^{(t+1)}$ using the following update step $\mathbf{x}^{(t+1)}\leftarrow\mathbf{x}^{(t)}+\eta\mathbf{v}^{(t)}$, where $\mathbf{v}^{(t)}=\operatornamewithlimits{argmax}_{\mathbf{w}\in\mathcal{P}}\langle\mathbf{w},\nabla\widehat{f}(\mathbf{x}^{(t)})\rangle$. For $\texttt{OPT}=\max_{\mathbf{x}\in\mathcal{P}}\widehat{f}(\mathbf{x})$ we have $(1-(1-\eta)^{T})\texttt{OPT}\leq\widehat{f}(\mathbf{x}^{(T)})+C\eta^{2}/2$. Here, the constant $C$ depends on the Lipschitz of the function, and $\mathbf{x}^{(T)}\in\mathcal{P}$ as it is a convex combination of vectors from the polytope. For $\eta=1/T$ and large enough $T$ we get $(1-1/\mathrm{e})\texttt{OPT}\leq\widehat{f}(\mathbf{x}^{(T)})+\epsilon$. Given $\mathbf{x}^{(T)}\in\mathcal{P}$, there are rounding procedures to obtain $S\in\mathcal{I}$ with $\widehat{f}(\mathbf{x}^{(T)})\leq f(S)$. The approximation factor $1-1/\mathrm{e}$ is the best possible assuming $\mathbf{P}\neq\mathbf{NP}$ (Feige, 1998).